Show calculations, and write clearly. Please submit via Blackboard. A) Calculate price of a share common stock based on the dividend discount model in which expected dividend-$1.35, Required return 10% and growth in earnings= 65%) 5Pts B) Calculate price to earnings multiple when expected earnings = $3.25 5 Pts 1. C) Calculate the PEG ratio 5Pts A) Calculate price to sales ratio when expected dividend net profit margin-20%, required return = 10 %, growth in earnings = 6.5% 2. 13 expected earnings= 3.25, 5Pts B) If expected sales are $100 per share what would be the fair value per share? 5 Pts
Answer #1

1a) The formulae is below:
price of share=expected dividend/(return-growth)
=1.35/(10%-6.5%)
=38.57
b)price to earnings= Market price/earning per share
=38.57/3.25
=11.87
c) PEG ratio= PE multiple/Earnings growth
=11.87/6.5
=1.83

2)Price to sales ratio=profit margin*payout*(1+growth)/(return-growth)
Payout= expected dividend/expected earning
=1.35/3.25
=41.54%
profit margin=20% growth=6.5% return=10%
=(20%*41.54%*(1+6.5%)/(10%-6.5%)
=2.53

b)Expected sales=100
Price to sales=2.53
fair value per share=expected sales*price to sales
=100*2.53
=253